(4y-4)^2+10(4y-4)+21=0 find all real and imaginary solutions to the equation
(4y-4)^2+40y-40+21=0
(4y-4)^2+40y+-19=0
(4y-4)^2+40y-19=0
0=0


Answer by dy5841(1)

4y-4)^2+40y-40+21=0
(4y-4)^2+40y+-19=0
(4y-4)^2+40y-19=0
0=0